OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

OpenStudy (anonymous):

still in segment 1 eh?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Yes, accursed thing. Thanks so much.

OpenStudy (anonymous):

@quietus @RadEn @trueseminole Still there?

OpenStudy (quietus):

hmm, sorry im still tryin lel

OpenStudy (anonymous):

yeah im just writing it down on paper

OpenStudy (anonymous):

so what i have so far is:

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x } + \frac{ 1 }{ x } = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

So by finding the LCD it will turn to \[\frac{ 3 }{ 3x } + \frac{ 3 }{ 3x } = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

long story short the answer is 4 but im trying to work it out on here

OpenStudy (anonymous):

\[\frac{ 3 }{ 3x }+\frac{ 3 }{ 3x } = \frac{ 2 }{ x}\]

OpenStudy (anonymous):

I have a question - wouldn't it be (x - 3) for Carl, not 3x? Because he doesn't take 3 times as long, he goes three times FASTER. I could be wrong, though.

OpenStudy (anonymous):

\[\frac{ 2 }{ x } =\frac{ 1 }{ 2 } = 4\]

OpenStudy (anonymous):

oh wow youre right it should be x + 3

OpenStudy (anonymous):

so let's start back over

OpenStudy (quietus):

Answer is 5

OpenStudy (quietus):